# Attention layers provably solve single-location regression

### v1

@@ Abstract.
Correspondence to pierre.marion@epfl.ch.
1. Introduction
@@
Attention-based models (Bahdanau et al., 2015), such as Transformer (Vaswani et al., 2017), have achieved unprecedented performance in various learning tasks, including natural language processing (NLP), e.g., text generation (Bubeck et al., 2023), translation (Luong et al., 2015), sentiment analysis (Song et al., 2019; Sun et al., 2019; Xu et al., 2019), and audio/speech analysis (Bahdanau et al., 2016). These developments have led to many architectural and algorithmic variants of attention-based models (see the review by Lin et al., 2022). At a high level, the success of attention has been linked to its ability to manage long-range dependencies in input sequences (Bahdanau et al., 2015; Vaswani et al., 2017), since attention consists in computing pairwise dependence between input tokens according to their projection in learned directions, independently of their location in the sequence.
@@
On the theoretical front, however, a deeper understanding of attention-based neural networks is still in its infancy. This limited progress is due both to the complexity of the architectures and to the disturbing diversity of relevant tasks. A common approach to tackle these challenges is to introduce a simplified task that models certain features of real-world tasks, followed by demonstrating a simplified version of the attention mechanism capable of solving the task. Prominent examples of this pattern include studying in-context learning with linearized attention (Ahn et al., 2023; von Oswald et al., 2023; Zhang et al., 2024), topic understanding with a single-layer attention and alternate minimization scheme (Li et al., 2023b), learning spatial structure with positional attention (Jelassi et al., 2022), and next-token prediction with certain bigram structures (Bietti et al., 2023; Tian et al., 2023).
While these works shed light on some abilities of Transformer, they do not encompass all the characteristics of tasks where Transformer performs well, in particular in NLP. Two features of particular interest, which to our knowledge have not been addressed in previous theoretical studies on Transformer, are token-wise sparsity, where relevant information is contained in a limited number of tokens, and internal linear representations, which are interpretable representations of the input constructed by the model.
@@
Contributions.  To understand why attention is a suitable architecture for addressing these features, we introduce single-location regression, a novel statistical task where attention-based predictors excel (Section 2). In a nutshell, this task is a regression problem with a sequence of tokens as input. The key novelty is that only one token determines the prediction, and the location of this token is a latent random variable that changes based on the input sequence. Consequently, solving the task requires first identifying the location of the relevant token, which can be done by learning a latent linear projection, followed by performing regression on that token.
@@
Organization.  Section 2 presents the mathematical framework of single-location regression, followed by motivations from language processing. Section 3 is dedicated to defining our predictor and explaining its connection with attention. We then move on to the mathematical study, from both statistical (Section 4) and optimization (Section 5) points of view. Section 6 presents the outline of the proof of our main mathematical result, while Section 7 concludes the paper.
2. Single-location regression task
@@
2.1. Statistical setting
We consider a regression scenario where the inputs are sequences of $L$ independent random tokens [^1] For the sake of simplicity, we interchangeably use the terms “token” and “embedding”, although they have different meanings in the NLP community. $(X_{1},\ldots,X_{L})$ taking values in $\mathbb{R}^{d}$. The output $Y\in\mathbb{R}$ is assumed to be given by
@@ 2.1. Statistical setting
A knee-jerk reaction would be to fit a linear model to the pair $(X_{1}^{\top},\ldots,X_{L}^{\top},Y)$. One might also consider tackling the problem with classical statistical approaches dedicated to sparsity, such as a Lasso estimator or a group-Lasso technique (Hastie et al., 2009). However, as we will see (in Section 4), all linear predictors fail due to the unknown and changing location of $J_{0}$. We note in addition that $\mathbb{E}[\|X_{\ell}\|_{2}^{2}]=d$ when $\ell\neq J_{0}$, while $\mathbb{E}[\|X_{J_{0}}\|_{2}^{2}]=d/2+\gamma^{2}d$. Therefore, choosing $\gamma^{2}=1/2$ implies that tokens are of the same squared norm in expectation, whether they are discriminatory of not. This shows that any approach based on comparing the magnitude of the tokens does not yield meaningful results. Ultimately, it is necessary to implement a more sophisticated approach, capable of taking into account the characteristics of the problem.
[…]
2.2. Language processing motivation
@@
(a) Examples of input-output pairs. The input is a text containing two sentences (e.g., a question and an answer), and the task is to perform sentiment analysis only for the second sentence. The $Y$ label is symbolized here by shades of color, where green (resp. red) corresponds to positive (resp. negative) feelings. The relevant information to determine the label is sparse, typically concentrated in a single token (in grey): changing this token flips the output label.
@@
(b) Accuracy of a logistic regression trained on embeddings of [CLS] tokens in the hidden layers of a pretrained Transformer model. Initial embeddings of [CLS] (at layer 00) are not context-aware, so they have a pure-chance accuracy of $50\%$. In hidden layers, the [CLS] token contains a representation of the sentence that is rich enough to achieve high scores with a linear classifier, and robust to out-of-distribution changes in token distribution and sentence structure.
@@
Figure 1. A simple sentiment analysis task with synthetic data, which exemplifies (a) token-wise sparsity and (b) internal linear representations. We refer to Appendix D for details on the experiment.
@@
Token-wise sparsity.  In many NLP tasks, the most relevant information is contained in a few tokens, where we recall that tokens correspond to small text units (typically, words or subwords), which are embedded in ${\mathbb{R}}^{d}$ using a learnt dictionary. This sparsity structure is revealed by the line of work on sparse attention (see, e.g., Martins and Astudillo, 2016; Niculae and Blondel, 2017; Correia et al., 2019; Child et al., 2019; Jaszczur et al., 2021; Kim et al., 2022; Farina et al., 2024), which obtains similar performance to full attention despite attending to a small fraction of tokens at each step. To illustrate this, we consider a simple sentiment analysis task in Figure 1(a), and observe that changing one token flips the output label. This is modeled in ($P_{\mathrm{learn}}$) by having the output $Y$ depend on a single token $J_{0}$, whose location furthermore varies depending on the input, as in NLP.
@@
Internal linear representations.  Linear projections of internal representations of Transformer (a.k.a. linear probing) allows to retrieve interpretable information about the input (see, e.g., Bolukbasi et al., 2021; Burns et al., 2023; Li et al., 2023a, and references therein). Such a linear structure is also present in the learned token embeddings that are fed as input to language models (Mikolov et al., 2013a, b; Bolukbasi et al., 2016; Nanda et al., 2023; Wen-Yi and Mimno, 2023). In our task ($P_{\mathrm{learn}}$), the two directions $k^{\star}$ and $v^{\star}$ have to be learned by the model in order to solve the task. Figure 2 gives an example of such directions for the toy task described above. Note that this illustration relies on initial embeddings, but we emphasize that similar structures also appear in the intermediate representations of Transformer. This is shown in Figure 1(b), where we observe that pretrained Transformer architectures indeed build internal representations that are sufficient to solve the toy NLP task with a linear classifier.
@@
Figure 2. Modeling of an NLP task within our statistical setting ($P_{\mathrm{learn}}$). The token embeddings $X_{1},\dots,X_{L}$ are constructed by adding the embeddings of each word and a positional encoding. For illustration purposes, we assume that each token corresponds to a word, and that the positional encoding solely depends on the part of the sentence (before or after the question mark), which differs from usual practice. Then, let the direction $k^{\star}$ encode both the notion of sentiment and the position in the second part of the sentence. Thus only the last token of the sentence is aligned (positively) with $k^{\star}$, and we have $J_{0}=L$. As for $v^{\star}$, it encodes whether the word is associated with a positive or negative sentiment. Note that several tokens are positively or negatively aligned with $v^{\star}$, but the output $Y$ only depends on the token $J_{0}$. This illustrates the interest of having two latent directions $k^{\star}$ and $v^{\star}$, one that filters the informative token and one that aligns with the output $Y$.
@@
We acknowledge that our modeling of NLP tasks presents limitations such as fixed sequence length, independent tokens, and output depending only on a single token. More complex and realistic models could be considered, but they require costly technical adjustments. Moreover, as we argue above, our problem ($P_{\mathrm{learn}}$) still preserves interesting aspects of NLP tasks, which makes it a relevant test bed for theoretical study of Transformer. Furthermore, from a statistical perspective, this is an original task that requires the implementation of a customized estimation strategy. It is precisely in this context that attention models prove their effectiveness, as we demonstrate next.
3. An attention-based predictor to solve the regression task
@@
Step 1: An oracle non-differentiable predictor.  If the vectors $(k^{\star},v^{\star})\in(\mathbb{S}^{d-1})^{2}$ were known, then a natural procedure to solve the task ($P_{\mathrm{learn}}$) would be to predict $Y$ from $\mathbb{X}$ via
@@
The $\operatorname*{arg\,max}$ part detects the location $J_{0}$ by exploiting the fact that all $X_{\ell}$ have zero mean except $X_{J_{0}}$, while the $\mathbb{X}v^{\star}$ part exploits the linear relationship $Y=X_{J_{0}}^{\top}v^{\star}+\xi$. In a more compact format, this ideal predictor can be rewritten as
$T(\mathbb{X})=\sum_{\ell=1}^{L}\mathbbb{1}_{\operatorname*{arg\,max}(\mathbb{X}k^{\star})=\ell}(\mathbb{X}v^{\star})_{\ell}\,,$
which is a linear regression in the direction $v^{\star}$ with non-differentiable weights depending on $k^{\star}$.
Step 2: A trainable predictor.  In practice, the vectors $k^{\star}$ and $v^{\star}$ are unknown and must be estimated from the data. In addition, the non-differentiability of the $\operatorname*{arg\,max}$ function poses significant optimization challenges. To solve this problem, the most common approach in machine learning is to replace $\operatorname*{arg\,max}$ with a softmax function with inverse temperature $\lambda>0$, i.e., for $z=(z_{1},\ldots,z_{L})\in~{}\mathbb{R}^{L}$, $[\mathrm{softmax}(\lambda z)]_{j}={e^{\lambda z_{j}}}/{\sum_{\ell=1}^{L}e^{\lambda z_{\ell}}}$. This leads us to the model
@@
where the $\mathrm{erf}$ function is applied component-wise. We emphasize that empirically, this simplification of softmax using a component-wise nonlinearity has been shown not to degrade performance in Transformer architectures (Hron et al., 2020; Shen et al., 2023; Wortsman et al., 2023).
Connection to attention.  It turns out that our estimation method finds a natural interpretation in terms of attention models. To see this, consider a model consisting of a single attention layer with a single head (Vaswani et al., 2017)
@@
where the dimensions $p,o\in\mathbb{N}^{*}$ are hyperparameters of the model, the softmax function is applied row by row, $Q,K,V\in\mathbb{R}^{d\times p}$ and $O\in\mathbb{R}^{o\times p}$ are the regular query, key, value, and output matrices, and $\lambda$ is usually taken to be $1/\sqrt{p}$. In practice, the attention head is added to $\mathbb{X}$ via a skip connection, which enforces $o=d$. In a nutshell, $K$ detects which tokens are relevant in the sentence, $V$ encodes the regression coefficient, and $Q$ encodes where we store the information.
@@
It is important to note that only considering the first output coordinate is a valid simplification for a single attention layer, but not when multiple layers are stacked, as all coordinates of the attention output contribute. Nevertheless, even in this latter more realistic case, the [CLS] token—or the similar concepts of attention sinks and registers—has been empirically shown to play a crucial role (Clark et al., 2019; Darcet et al., 2024; Xiao et al., 2024). This is also confirmed by our experiment in Figure 1(b), where we show that the [CLS] token in pretrained Transformer architectures stores an internal representation of the sentence that is sufficient to solve simple NLP tasks with a linear classifier. This further motivates the need to understand how information is stored in this token.
It turns out that there is a direct connection between the model $T_{\lambda}^{(\text{soft},k,v)}(\mathbb{X})$ defined in (2) and the attention model $T_{\lambda}^{(Q,K,V,O)}(\mathbb{X})_{1}$ described in (5). To see this, take $o=1$, to adapt the model (5) for univariate regression, and set $p=1$, a reasonable assumption given both empirical and theoretical evidence suggesting that Transformer parameter matrices are low-rank (Aghajanyan et al., 2021; Kajitsuka and Sato, 2024). Then, let $Q\in{\mathbb{R}}^{d\times 1}$ be any vector with positive correlation with $X_{\textnormal{[CLS]}}$ (for instance it suffices to take $Q=X_{\textnormal{[CLS]}}$), and $O=1$. We then deduce that
@@
In other words, the attention layer (5) matches the considered predictor in (2) with a softmax inverse temperature proportional to the scalar product between $X_{\textnormal{[CLS]}}$ and $Q$. Thus, our results, in particular the study of the training dynamics in Section 5, can be seen as a model of how Transformer builds internal linear representations of the input during training.
4. Risk of the oracle and of the linear predictors

### v2

@@ Abstract
1 Introduction
@@
Attention-based models (Bahdanau et al., 2015), such as Transformer (Vaswani et al., 2017), have achieved unprecedented performance in various learning tasks, including natural language processing (NLP), e.g., text generation (Bubeck et al., 2023), translation (Luong et al., 2015), sentiment analysis (Song et al., 2019; Sun et al., 2019; Xu et al., 2019), and audio/speech analysis (Bahdanau et al., 2016). These developments have led to many architectural and algorithmic variants of attention-based models (see the review by Lin et al., 2022). At a high level, the success of attention has been linked to its ability to manage long-range dependencies in input sequences (Bahdanau et al., 2015; Vaswani et al., 2017), since it consists in computing pairwise dependence between input tokens according to their projection in learned directions, independently of their location in the sequence.
@@
On the theoretical front, however, a deeper understanding of attention-based neural networks is still in its infancy. This limited progress is due both to the complexity of the architectures and to the disturbing diversity of relevant tasks. A common approach to tackle these challenges is to introduce a simplified task that models certain features of real-world tasks, followed by demonstrating a simplified version of the attention mechanism capable of solving the task. Prominent examples of this pattern include studying in-context learning with linearized attention (Ahn et al., 2023; von Oswald et al., 2023; Zhang et al., 2024), topic understanding with single-layer attention and alternate minimization scheme (Li et al., 2023b), learning spatial structure with positional attention (Jelassi et al., 2022), next-token prediction with latent bigram (Bietti et al., 2023; Tian et al., 2023) or causal graph (Nichani et al., 2024) structures, and sparse token selection (Wang et al., 2024). We refer to Appendix F for additional discussion on some of these related works. While these works shed light on some abilities of Transformer, they do not encompass all the characteristics of tasks where Transformer performs well, in particular in NLP. Two features of particular interest, which to our knowledge have not been addressed in previous theoretical studies on Transformer, are token-wise sparsity, where relevant information is contained in a limited number of tokens, and internal linear representations, which are interpretable representations of the input constructed by the model.
Contributions.
@@
To understand why attention is a suitable architecture for addressing these features, we introduce single-location regression, a novel statistical task where attention-based predictors excel (Section 2). In a nutshell, this task is a regression problem with a sequence of tokens as input. The key novelty is that only one token determines the prediction, and the location of this token is a latent random variable that changes based on the input sequence. Consequently, solving the task requires first identifying the location of the relevant token, which can be done by learning a latent linear projection, followed by performing regression on that token.
@@
Organization.
Section 2 presents the mathematical framework of single-location regression, followed by motivations from language processing. Section 3 is dedicated to defining our predictor and explaining its connection with attention. We then move on to the mathematical study, from both statistical (Section 4) and optimization (Section 5) points of view. Section 6 concludes the paper.
2 Single-location regression task
@@
2.1 Statistical setting
We consider a regression scenario where the inputs are sequences of $L$ random tokens [^1] For the sake of simplicity, we interchangeably use the terms “token” and “embedding”, although they have slightly different meanings in the NLP community. $(X_{1},\ldots,X_{L})$ taking values in $\mathbb{R}^{d}$. The output $Y\in\mathbb{R}$ is assumed to be given by
@@ 2.1 Statistical setting
A knee-jerk reaction would be to fit a linear model to the pair $(X_{1}^{\top},\ldots,X_{L}^{\top},Y)$. One might also consider tackling the problem with classical statistical approaches dedicated to sparsity, such as a Lasso estimator or a group-Lasso technique (Hastie et al., 2009). However, as we will see (in Section 4), all linear predictors fail due to the unknown and changing location of $J_{0}$. We note in addition that $\mathbb{E}[\|X_{\ell}\|_{2}^{2}]=d$ when $\ell\neq J_{0}$, while $\mathbb{E}[\|X_{J_{0}}\|_{2}^{2}]=d/2+\gamma^{2}d$. Therefore, choosing $\gamma^{2}=1/2$ implies that tokens have the same squared norm in expectation, whether they are discriminatory of not. This shows that any approach based on comparing the magnitude of the tokens does not yield meaningful results. Ultimately, it is necessary to implement a more sophisticated approach, capable of taking into account the characteristics of the problem.
2.2 Language processing motivation
@@
(a) Examples of input-output pairs. The input is a text containing two sentences (e.g., a question and an answer), and the task is to perform sentiment analysis only for the second sentence. The $Y$ output is symbolized here by a color code, where green (resp. red) corresponds to positive (resp. negative) feelings. The relevant information is sparse, typically concentrated in a single token: changing the grey token flips the output.
@@
(b) Accuracy of logistic regression on embeddings of [CLS] tokens in the hidden layers of a pretrained Transformer model. Initial embeddings of [CLS] (at layer 00) are not context-aware, so they have a pure-chance accuracy of $50\%$. In hidden layers, the [CLS] token contains a representation of the sentence that achieves high scores and is robust to out-of-distribution changes in token distribution and sentence structure.
@@
Figure 1: A simple sentiment analysis task with synthetic data, which exemplifies (a) token-wise sparsity and (b) internal linear representations. We refer to Appendix E for details on the experiment.
Token-wise sparsity.
@@
In language tasks, the relevant information is often contained in few tokens, where we recall that tokens correspond to small text units (typically, words or subwords), which are embedded in ${\mathbb{R}}^{d}$ using a learned dictionary. This sparsity is revealed by the success of sparse attention (Martins & Astudillo, 2016; Niculae & Blondel, 2017; Correia et al., 2019; Child et al., 2019; Jaszczur et al., 2021; Kim et al., 2022; Farina et al., 2024), which is competitive with full attention while attending to fewer tokens. As an illustration, we consider a simple sentiment analysis task in Figure 1(a), and observe that changing one token flips the output. This is modeled in ($P_{\mathrm{learn}}$) by having the output $Y$ depend on a single token $J_{0}$, whose location furthermore varies with the input.
Internal linear representations.
@@
Linear projections of internal representations of Transformer (a.k.a. linear probing) contain interpretable information (Bolukbasi et al., 2021; Burns et al., 2023; Li et al., 2023a). Such a linear structure is also present in the learned token embeddings that are fed as input to language models (Mikolov et al., 2013a; b; Bolukbasi et al., 2016; Nanda et al., 2023; Wen-Yi & Mimno, 2023). In our task ($P_{\mathrm{learn}}$), the two directions $k^{\star}$ and $v^{\star}$ have to be learned by the model in order to solve the task. Figure 2 gives an example of possible such directions for the toy task described above. While this illustration relies on initial embeddings, similar structures also appear in the intermediate representations of Transformer. This is shown in Figure 1(b), where we observe that pretrained Transformer architectures indeed build internal representations that are sufficient to solve the task with a linear classifier.
@@
Figure 2: Modeling of an NLP task within our statistical setting ($P_{\mathrm{learn}}$). The token embeddings $X_{1},\dots,X_{L}$ are constructed by adding the embeddings of each word and a positional encoding. For illustration purposes, we assume that each token corresponds to a word, and that the positional encoding solely depends on the part of the sentence (before or after the question mark), which differs from usual practice. Then, let the direction $k^{\star}$ encode both the notion of sentiment and the position in the second part of the sentence. Thus only the last token of the sentence is aligned (positively) with $k^{\star}$, and we have $J_{0}=L$. As for $v^{\star}$, it encodes whether the word is associated with a positive or negative sentiment. Note that several tokens are positively or negatively aligned with $v^{\star}$, but the output $Y$ only depends on the token $J_{0}$. This illustrates the interest of having two latent directions $k^{\star}$ and $v^{\star}$, one that filters the informative token and one that aligns with the output $Y$.
@@
We acknowledge that our statistical task presents limitations such as fixed sequence length, independent tokens, and output depending only on a single token. More complex models could be considered, but at significant technical cost. Moreover, as argued above, our problem ($P_{\mathrm{learn}}$) preserves interesting aspects of NLP tasks, which makes it relevant for theoretical study of Transformer. Furthermore, it is an original statistical task requiring the implementation of a customized estimation strategy. It is precisely in this context that attention models prove their effectiveness, as we show next.
3 An attention-based predictor to solve the regression task
@@
Step 1: An oracle non-differentiable predictor.
If the vectors $(k^{\star},v^{\star})\in(\mathbb{S}^{d-1})^{2}$ were known, then a natural procedure to solve the task ($P_{\mathrm{learn}}$) would be to predict $Y$ from $\mathbb{X}$ via
@@
The $\operatorname*{arg\,max}$ part detects the location $J_{0}$ by exploiting the fact that all $X_{\ell}$ have zero mean except $X_{J_{0}}$, while the $\mathbb{X}v^{\star}$ part exploits the linear relationship $Y=X_{J_{0}}^{\top}v^{\star}+\xi$. In a more compact format, this ideal predictor can be rewritten as $T(\mathbb{X})=\sum_{\ell=1}^{L}\mathbbb{1}_{\operatorname*{arg\,max}(\mathbb{X}k^{\star})=\ell}(\mathbb{X}v^{\star})_{\ell}\,,$ which is a linear regression in the direction $v^{\star}$ with non-differentiable weights depending on $k^{\star}$.
Step 2: A trainable predictor.
In practice, the vectors $k^{\star}$ and $v^{\star}$ are unknown and must be estimated from the data. In addition, the non-differentiability of the $\operatorname*{arg\,max}$ function poses significant optimization challenges. To solve this problem, the most common approach in machine learning is to replace $\operatorname*{arg\,max}$ with a softmax function with inverse temperature $\lambda>0$, i.e., for $z=(z_{1},\ldots,z_{L})\in~{}\mathbb{R}^{L}$, $[\mathrm{softmax}(\lambda z)]_{j}={e^{\lambda z_{j}}}/{\sum_{\ell=1}^{L}e^{\lambda z_{\ell}}}$. This leads us to the model
@@
where the $\mathrm{erf}$ function is applied component-wise. The choice of this activation function enables closed-form expectations for functions of Gaussian random variables (see, e.g., Lemma 18). Note that the role of softmax in attention is an open question in the community. Several empirical papers investigate simplifying softmax into a component-wise nonlinearity (Qin et al., 2022; Shen et al., 2023; Wortsman et al., 2023; Ramapuram et al., 2024), and have observed a similar performance. These works emphasize the importance of the normalization $\lambda$ when replacing softmax, which we also find out to play an important role (see Corollary 2 and Section 5).
Connection to attention.
It turns out that our estimation method finds a natural interpretation in terms of attention models. To see this, consider a model consisting of a single attention layer with a single head (Vaswani et al., 2017)
@@
where the dimensions $p,o\in\mathbb{N}^{*}$ are hyperparameters of the model, the softmax function is applied row by row, $Q,K,V\in\mathbb{R}^{d\times p}$ and $O\in\mathbb{R}^{o\times p}$ are the regular query, key, value, and output matrices, and $\lambda$ is usually taken to be $1/\sqrt{p}$. In practice, the attention head is added to $\mathbb{X}$ via a skip connection, which enforces $o=d$. In a nutshell, $K$ detects which tokens are relevant in the sentence, $V$ encodes the regression coefficient, and $Q$ encodes where to store the information.
@@
It is important to note that only considering the first output coordinate is a mathematically valid simplification for a single attention layer, but not when multiple layers are stacked, as all coordinates of the attention output contribute. Nevertheless, even in this latter more realistic case, the [CLS] token—or the similar concepts of attention sinks and registers—has been empirically shown to play a crucial role (Clark et al., 2019; Darcet et al., 2024; Xiao et al., 2024). This is also confirmed by our experiment in Figure 1(b), where we show that the [CLS] token in pretrained Transformer architectures stores an internal representation of the sentence that is sufficient to solve simple NLP tasks with a linear classifier. This further motivates the need to understand how information is stored in this token.
It turns out that there is a direct connection between the model $T_{\lambda}^{(\text{soft},k,v)}(\mathbb{X})$ defined in (2) and the attention model $T_{\lambda}^{(Q,K,V,O)}(\mathbb{X})_{1}$ described in (5). To see this, take $o=1$, to adapt the model (5) for univariate regression, and set $p=1$, a reasonable assumption given both empirical and theoretical evidence suggesting that Transformer parameter matrices are low-rank (Aghajanyan et al., 2021; Kajitsuka & Sato, 2024). Then, let $Q\in{\mathbb{R}}^{d\times 1}$ be any vector with positive correlation with $X_{\textnormal{[CLS]}}$ (for instance it suffices to take $Q=X_{\textnormal{[CLS]}}$), and $O=1$. We then deduce that
@@
In other words, the attention layer (5) matches the considered predictor in (2) with a softmax inverse temperature proportional to the scalar product between $X_{\textnormal{[CLS]}}$ and $Q$. Thus, our results, in particular the study of the training dynamics in Section 5, can be seen as a model of how Transformer builds internal representations of the input during training. This is also supported by numerical experiments showing that Transformer layers behave similarly to our predictor (see Appendix E).
4 Risk of the oracle and of the linear predictors